OpenStudy (anonymous):

At the Stop 'n Go tune-up and brake shop, the manager has found that an SUV will require a tune-up with a probability of 0.6, a brake job with a probability of 0.1 and both with a probability of 0.02. What is the probability that an SUV requires either a tune up or a brake job

OpenStudy (anonymous):

nevermind got the answer

OpenStudy (phoenixfire):

From "either a tune-up or a brake job" I figure it means one or the other occurs, but not both. In other words an exclusive or \[P(A) + P(B) - 2P(A)P(B)=0.6+0.1-(2)(0.6)(0.1)=0.7-0.12=0.58\]

OpenStudy (phoenixfire):

You got the answer... Was I right?

OpenStudy (anonymous):

I got 0.68

OpenStudy (anonymous):

0.6 + 0.1 - 0.02 0.7 - 0.02 = 0.68

OpenStudy (phoenixfire):

Do you have the correct answers? Is 0.68 correct?

OpenStudy (phoenixfire):

Ah I see. 0.68 is correct \[P(A)+P(B)-P(A and B)\]

OpenStudy (anonymous):

no i didnt have the correct answers but I found out the equation to it